Given that purchasing power parity holds, what would be the price of oil in terms of Peruvian sols. Use the following information for your calculations:

Gold = $300 per ounce in U.S.

Gold = 2,400 sols per ounce in Peru

Oil = $25 per barrel in U.S.

A. 3.125

B. 96

C. 250

D. 200



1
Expert's answer
2021-12-10T13:31:02-0500

"Purchasing power Parity= \frac{Cost of good X in country 1}{Cost of good X in country 2}"

"Purchasing power Parity= \frac{Cost of gold in US}{Cost of gold in Peru}=\frac{300}{2,400}=0.125"

When purchasing power parity is held, then the price of oil in terms of Peruvian sols will be;

"=\frac{25}{Cost of oil in Peru}=0.125"

"Cost of oil in Peru=\frac{25}{0.125}=200Peruvian sols"


"=200 Peruvian sols"